OpenStudy (anonymous):

x2 -7x + 10 = 0 HELP MEEE i need solution set

OpenStudy (mimi_x3):

\[a+b = -5 + -2\] \[a*b = -5*-2\] So.. \[(x-5)(x-2)\] \[x = 5,2\]

OpenStudy (anonymous):

(x-5)(x-2) = 0 x = 5 or 2